OpenStudy (anonymous):

Evaluate the Integral: 3/x from 4 to 2. I'm really bad at integrals, I totally forgot how to do them, please explain how to do it.

OpenStudy (anonymous):

3/x (dx)

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

\[\int_2^4\frac{3}{x}dx\] Now first of all, do you know the integral of \(\large \frac{1}{x}\)?

OpenStudy (anonymous):

isnt that ln(x)

OpenStudy (anonymous):

Yes, do you remember the fundamental theorem of calculus?

OpenStudy (anonymous):

we dont go over theorems in class :/

OpenStudy (anonymous):

\[\large \int_2^4\frac{3}{x}dx = 3 \int_2^4\frac{1}{x}dx = 3\ln(x) \Big]_2^4\]

OpenStudy (anonymous):

So 3(ln(4)-ln(2)) = ?

OpenStudy (anonymous):

The fundamental theorem of calculus tells us to do a definite integral, we take the anti derivative (in this case \(3\ln(x)\)) and

OpenStudy (anonymous):

Yeah you got it

OpenStudy (anonymous):

okay :) 2.07 is the final answer. Tyvm. I will try to do the next problems and if anything, I will post a new thread.